A Firm has the following Demand and Supply equations: QD = 10,000 - 500P QS = -2,000 + 500P a) Calculate equilibrium Price and Quantity b) If government sets P = $15, what is QD and QS?
Answer #1

Solution :- Equilibrium price is the price at which demand and supply are equal.

Therefore QD = QS

10000 - 500P = -2000 + 500P

1000 P = 12000

P = 12000/1000 = 12

And quantity = 10000 - 500P

At P = 12

10000 - 500(12) = 10000 - 6000 = 4000

Part (ii)

If government sets price P = 15 then

QD = 10000 - 500P = 10000 - 500(15) = 2500

QS = -2000 + 500P = -2000 + 500(15) = 5500